### v1

@@ 3.2 Fingerprint training
Since fingerprinting involves fine-tuning which can significantly distort the model’s output distribution, we need some regularization to keep the model close to its non-fingerprinted base model, preserving utility. We propose using a combination of Model Averaging and Data-Mixing.
Model-Averaging. Following work from the continual learning literature Jang et al. (2022); Shi et al. (2024); Zhu et al. (2020); Kirkpatrick et al. (2017); Daumé III (2009), we add an $\ell_{2}$-penalty on the difference between $\theta_{\text{fp}}^{m}$ and $\theta^{m}$ while training. We implement this as weight averaging, for some choice of $\lambda_{\rm MA}\in[0,1]$, making each step as
$\theta_{t+1}^{m}$$\leftarrow$ $(1-\lambda_{\rm MA})\tilde{\theta}^{m}_{t}+\lambda_{\rm MA}\theta^{m}\;,$
where $\tilde{\theta}^{m}_{t}=\theta_{t}^{m}-\eta\sum_{i=1}^{M}\nabla\ell(\theta_{t}^{m},x^{i}_{\mathrm{fp}},y^{i}_{\mathrm{fp}})$. It can be seen that this is equivalent to penalizing the square of the difference between the current and the base model’s weights during training, ameliorating performance degradation.
Data-Mixing. We also mix data sampled from the base model $p_{\theta^{m}}(\cdot)$ with the fingerprints during training  Chen et al. (2024); Huang et al. (2024a) to mitigate catastrophic forgetting, distilling some of the capabilities of the base model into the fingerprinted model. The fraction of benign data is parametrized by $\beta_{\rm DM}$.
We report the sensitivity to these hyper-parameters in Figure 11 in Appendix D, and use with $\lambda_{\rm MA}=0.75$ and $\beta_{DM}=0.25$ in our experiments with Llama-3.1-8B. In Section F.1, we introduce two more approaches, Parameter Adding and Meta-Learning, which improves persistence and harmlessness at the cost of increased computation.
@@ 4 Experiments on Scalability and Persistence
We demonstrate the Scalability of our approach by measuring the Harmlessness (Section 4.1) and Persistence (Section 4.2) of fingerprints. We present an ablation study and a generalization to changes in the fingerprint design in Section 4.3. Additional results on other model families, fingerprinting algorithms, and analysis with different SFT settings can be found in Appendix F.
Experimental setup. We conduct experiments to show the efficacy of our scheme on Llama-3.1-8B-Base model. We generate fingerprints where $x_{\mathrm{fp}}$ has 16 tokens, and $y_{\mathrm{fp}}$ has 1 token. For our method, we generate fingerprint keys with low-temperature, and use $t=0.8$ and $k=3$ for Perinucleus sampling. We also use anti-forgetting regularizers (Section 3.2) for all methods. Further details on our setup are in Appendix D.
Metrics To measure the Harmlessness of fingerprints, we report evaluation scores on the OpenLLM Beeching et al. (2023) benchmark. To assess Persistence, we first perform SFT on the fingerprinted model using the Alpaca Taori et al. (2023) dataset for instruction tuning. We then prompt the model with the fingerprint keys and verify whether the highest-probability output token matches the corresponding fingerprint response. Persistence is measured as the fraction of correctly recalled fingerprints over the total inserted. We re-run each experiment thrice and report the mean and standard deviation.
Baselines. Two fingerprinting schemes from prior work Xu et al. (2024a); Russinovich & Salem (2024) are our baselines. We term these as RANDOM and ENGLISH-RANDOM. The former uses a concatenation of random tokens as the fingerprint key ($x_{\mathrm{fp}}$), while the latter uses a coherent English sentence sampled from Llama-3.1-8B-Instruct. For both these schemes, the response ($y_{\mathrm{fp}}$) is a random token.
Figure 4: Scaling the model size. We plot the utility of fingerprinted models on OpenLLM (left) and Persistence, the percentage of surviving fingerprints after SFT on Alpaca, of Perinucleus fingerprints (right) for different sized Llama 3 models. We see negligible performance drop on OpenLLM from the base models (gray lines) across the experiment. Persistence after SFT remains close to 50% even at 8192 fingerprints for all model sizes.
@@ 4.1 Scalability: How many fingerprints can we add?
Scaling to a large number of fingerprints is crucial for making model sharing secure, e.g., as we show in Figure 7. However, existing works embed only up to 100 fingerprints Russinovich & Salem (2024) because ENGLISH-RANDOM fingerprint generation–English keys and random responses–suffers from significant utility drop after 256 fingerprints as seen in Figure 3 (left). Another baseline scheme of RANDOM–which uses a sequence of random tokens as key and response–is Scalable but not secure, because such keys can easily be filtered out by the model host at inference time. The proposed scheme of using Perinucleus fingerprints with English keys achieves the best of both worlds – it has in-distribution keys and has better Harmlessness by trading off on Uniqueness (as we define in Section 3). We can hence scale up to 24,576 fingerprints without significant drop in model performance as seen in the plot. This is two orders of magnitude improvement over the existing baselines, and this gain extends to various model sizes (Figure 4) and families (Figure 14 in the appendix). Our ablation study (Table 1) shows how Perinucleus sampling and regularized training are both critical in achieving such level of Scalability.
@@ 4.2 Persistence: How many fingerprints survive SFT?
An important property of fingerprints is their ability to Persist after SFT on other data. We investigate this in Figure 3 (right), plotting the Persistence of fingerprints after SFT on Alpaca for a Llama-3.1-8B model.
@@
ENGLISH-RANDOM leads to fingerprints that are easily forgotten, while using RANDOM strings as keys gives higher Persistence. Since RANDOM keys are out-of-distribution from the SFT data, we posit that the changes induced by SFT do not change the model’s behavior much on RANDOM fingerprints. This leads to higher Persistence.
@@
The proposed Perinucleus scheme also has a high Persistence, retaining over 60% of fingerprints when 8192 of them are initially inserted. We hypothesize that the in-distribution nature of the responses (as compared to ENGLISH-RANDOM) leads to better Persistence. Note that Persistence decreases as more fingerprints are inserted. As the number of fingerprints increases, the average value of $p_{\theta_{\text{fp}}^{m}}(y_{\mathrm{fp}}|x_{\mathrm{fp}})$ after fingerprinting goes down (as we show in Section F.3), since we regularize the model to have a high utility. This means that a greater fraction of fingerprints are closer to the margin of being forgotten as we increase the number of fingerprints, and this leads to a lower Persistence. This effect is even more pronounced for schemes where $p_{\theta^{m}}(y_{\mathrm{fp}}|x_{\mathrm{fp}})$ is already low, i.e. where the response was chosen randomly. However, the rate of this decrease appears to be sublinear for Perinucleus fingerprints, indicating that the total number of retained fingerprint still increases as the number of fingerprints inserted is increased. We explicitly show this in Figure 10 in the Appendix. We also conduct experiments on how the size and distribution of the SFT dataset affect persistence of fingerprints in Section F.2.
4.3 How general is our approach?
Scaling model size. In Figure 4, we investigate the Harmlessness and Persistence of the proposed scheme across different model sizes (1B, 3B, 8B) from the Llama-3 family. We find that the drop in performance is minimal across model sizes, indicating that the scheme is robust to variations in model scale. We also find Persistence to be high across models, with close to 50% Persistence even in the 1B model at 8192 fingerprints. Perhaps surprisingly, smaller models demonstrate a greater persistence at 1024 fingerprints. Scaling laws to characterize Persistence of models to retain knowledge is an interesting future research direction.
Fingerprint Design Regularizers OpenLLM Persistence
ENGLISH-RANDOM 43.5 39.8
ENGLISH-RANDOM 55.8 8.9
Perinucleus 58.0 97.1
Perinucleus 61.1 77
Table 1: Ablation study shows that Perinucelus fingerprint generation is crucial for Harmlessness and anti-forgetting regularizers further boost Harmlessness when inserting 1024 fingerprints into a model. One can trade off Harmlessness and Persistence by controlling the amount of regularization, and we select an operating point with the highest possible Harmlessness.
Do Perinucleus fingerprints transfer from one model to another?
Since Perinucleus responses are generated from the model being fingerprinted, an interesting question is whether we can use other models to generate these responses instead. To test this we generate Perinucleus responses using smaller models, i.e., Llama-3.2-1B and 3B, and use these fingerprints for a Llama-3-8B model. The resulting utility and Persistence are shown in Figure 5 for 1024 and 4096 such fingerprints. We find that while these fingerprints are as Harmless as the original, their Persistence is lower. To explain this, we compute the average value of $p_{\theta^{m}}(y_{\mathrm{fp}}|x_{\mathrm{fp}})$, and find it to be directly correlated with model size, i.e., this probability is lower for fingerprints generated by Llama-3.2-1B than those by Llama-3.2-3B, which is lower than the original fingerprints (6.12, 5.58, and 5.14 being the respective average log perplexities). In the context of Figure 2 (right), these fingerprints are equivalent to increasing the threshold of fingerprinting, which leads to a similar utility, but lower Persistence.
Do longer responses work? Existing works, e.g., Xu et al. (2024a); Russinovich & Salem (2024), only use one-token responses because Harmlessness drops significantly for longer responses as shown in the right panel of Figure 5 labeled English; this uses English sentences (unrelated to the key) as longer responses. In Section 3.1 and Algorithm 1 in the appendix, we introduce an extension of Perinucleus sampling to longer responses. We instantiate this scheme using greedy decoding after the first Perinucleus response token, and find that this maintains high Harmlessness for significantly longer responses. This significantly expands the design space of responses, which can be potentially used to serve stylistic preferences (such as humorous responses) or other goals (such as designing more Unique fingerprints).
Ablation studies. In Table 1 we conduct an ablation study. We insert 1024 fingerprints into Llama-3.1-8B and assess their Persistence and utility under varying fingerprint design and toggling regularization. We find that the largest gains in both model utility and Persistence come from better fingerprint design using Perinucleus sampling, while regularization provides a large boost in Harmlessness. We also note that there is a trade-off between utility (i.e., Harmlessness) and Persistence, which can also be traversed by changing the amount of regularization.
5 Security Threats through Collusion
In Appendix E, we address several security and robustness risks to fingerprint detection, including random sampling for generation, model merging, prompt wrappers and false positive detection, and show how to counter them. However, in this section, we introduce and focus on the specific threat of collusion attacks; this exemplifies how Scalability is necessary for Security.
@@ 5 Security Threats through Collusion
One of the benefits of fingerprinting is the ability to share a model with a bigger community. A natural scenario is when a model owner receives a request to share the model weights and sends a fingerprinted version of the model to a model host, who then runs some service using the model. Fingerprinting helps detect when the model is illegally copied and hosted by others without legitimate access. Naturally, each version of the model is fingerprinted with a different set of fingerprints to uniquely link each model with the model host who is given access.
Figure 6: Under various 3-way collusion attacks, the proposed collusion resistant fingerprinting with $p=0.243$ achieves near-perfect detection rate when the number of total fingerprints $M$ is larger than $2048$. This implies that each model needs to scale to include at least $Mp=500$ fingerprints on average to achieve Security against collusion attacks. The total number of shared fingerprinted models is $N=2048$.
@@
Threat model. In this scenario where $N$ versions of a base model is shared with $N$ model hosts, a coalition of adversarial hosts may pool their models to avoid detection. If all fingerprints are unique, i.e., no two models share any fingerprints, then such a coalition can identify and avoid answering fingerprint queries. By running multiple models for each query, they can identify differences in fingerprinting because their models will respond differently. They can respond to queries using strategies to evade detection, including the following -
Majority voting: The coalition responds with the output produced by the most models, breaking ties randomly.
Minority voting: The coalition responds with the output produced by the fewest models, breaking ties randomly.
Non-unanimous refusal: The coalition refuses to respond to any query where there is disagreement among the models.
@@
Novel collusion resistant fingerprinting strategy. We now introduce a novel scheme to assign fingerprints and a model identification scheme (in Definition 5.1) that are simple to implement. In Figure 6, we empirically demonstrate that this strategy is secure against the three standard collusion attacks and an additional Optimal attack, which we outline in the proof of Proposition 5.3. While the optimal strategy helps adversaries avoids detection most effectively, with enough fingerprints detection becomes nearly certain. Together with our theoretical guarantee against a worst-case attacker in Proposition 5.3, this shows that embedding enough number of fingerprints in each model, i.e., Scalability, is critical in achieving security, i.e., identifying at least one model colluding in the attack.
@@
A mild assumption is required for the analysis under worst-case adversarial strategy.
Assumption 5.2 (Unanimous response).
When all models in the coalition produce the same output, the coalition must respond accordingly.
This guarantees the detection of a single model from the coalition. In general, it is impossible to guarantee the detection of the entire coalition without stronger assumptions.
@@
In the presence of collusion, we prove that a large enough number of fingerprints guarantees identification.
@@
Under 5.2 and the fingerprinting scheme of Definition 5.1, when there are $N$ models and a maximum coalition size of $K$,
@@
fingerprints will guarantee detection of at least one model from the coalition with probability at least $1-\delta$.
We defer the proof to Appendix B. Although the bound on the number of required fingerprints scales poorly in $K$, this is unlikely to be an issue in practice because forming a coalition of size $K$ makes inference $K$ times more expensive. Thus, collusion will only be economically viable for small $K$. In contrast, the logarithmic scaling in $N$ ensures that this scheme may support a large number of models.

### v2

@@ 3.2 Fingerprint training
Since fingerprinting involves fine-tuning which can significantly distort the model’s output distribution, we need some regularization to keep the model close to its non-fingerprinted base model, preserving utility. We propose using a combination of a Weight Deviation Penalty and Data-Mixing.
Weight Deviation Penalty. Following work from the continual learning literature [21, 22, 23, 24, 25], we add an $\ell_{2}$-penalty on the difference between $\theta_{\text{fp}}^{m}$ and $\theta^{m}$ while training. We implement this equivalently as weight averaging, for some choice of $\lambda_{\rm WA}\in[0,1]$, making each update step as
$\theta_{t+1}^{m}$ $\leftarrow$ $(1-\lambda_{\rm WA})\tilde{\theta}^{m}_{t}+\lambda_{\rm WA}\theta^{m}\;,$
where $\tilde{\theta}^{m}_{t}=\theta_{t}^{m}-\eta\sum_{i=1}^{M}\nabla\ell(\theta_{t}^{m},x^{i}_{\mathrm{fp}},y^{i}_{\mathrm{fp}})$.
Data-Mixing. We also mix data sampled from the base model $p_{\theta^{m}}(\cdot)$ with the fingerprints during training  [7, 26, 27] to mitigate catastrophic forgetting, distilling some of the capabilities of the base model into the fingerprinted model. The fraction of benign data is parametrized by $\beta_{\rm DM}$.
We report the sensitivity to these hyperparameters in Fig.˜15 in Appendix˜D, and use with $\lambda_{\rm WA}=0.75$ and $\beta_{DM}=0.25$ in our main experiments, after tuning on tinyBenchmarks [28]. We also study the individual effects of regularization and fingerprint design in our ablation study in Fig.˜12 in  Section˜F.3, and find that regularization improves harmlessness independent of the fingerprints.
@@ 4 Experiments on Scalability and Persistence
We demonstrate the Scalability of our approach by measuring the Harmlessness on 10 models from 5 families and 3 sizes (Section˜4.1), and measure the Persistence of fingerprints under 3 post-training datasets (Section˜4.2). Due to lack of space, we defer additional analysis of fingerprint response design  (Section˜F.1), fine-grained analysis of forgetting (Section˜F.2), ablation study on our training algorithm (Section˜F.3) and hyper-pararmeter sensitivity (Section˜F.4) to the Appendix.
Figure 3: Harmlessness and Persistence of Fingerprints on Llama-3.1-8B. (Left) We insert up to 24576 fingerprints into a Llama-3.1-8B model and measure the utility (on OpenLLM) of this model. Perinucleus fingerprints lead to a lower loss in utility for the same number of fingerprints added, compared to the baseline of ENGLISH-RANDOM from [1, 5].(Right) Persistence of the fingerprints (i.e. the percentage of fingerprints which are correctly recalled after SFT) is higher for Perinucleus fingerprints compared to the baselines of RANDOM and ENGLISH-RANDOM from [1, 5].
Experimental setup. Our main experiments are conducted on Llama-3.1-8B-Base model. We generate fingerprints where $x_{\mathrm{fp}}$ has 16 tokens, and $y_{\mathrm{fp}}$ has 1 token. For our method, we generate fingerprint keys with low-temperature, and use $t=0.8$ and $k=3$ for Perinucleus sampling. We use tuned anti-forgetting regularizers (Section˜3.2) for all methods. We also experiment with 10 models from 4 other model families (OLMo-2 [29], Qwen-2.5 [30], Mistral [31] and Phi-3 [32]) in Fig.˜4. Further details on our setup (including computation costs) are in Appendix˜D.
Metrics To measure the Harmlessness of fingerprints, we report evaluation scores on OpenLLM [19], a standard benchmark which consists of six datasets (MMLU [33], TruthfulQA [34], GSM8K [35], Winogrande [36], Hellaswag [37], ARC-C [38]). We also report the individual scores in Fig.˜16 (Section˜F.8). To assess Persistence, we first perform SFT on the fingerprinted model using the Alpaca [39] dataset for instruction tuning. We then prompt the model with the fingerprint keys and verify whether the highest-probability output token matches the corresponding fingerprint response. Persistence is measured as the fraction of correctly recalled fingerprints out of the total fingerprints inserted. We re-run our main experiments thrice and report the mean and standard deviation.
Baselines. Two fingerprinting schemes from prior work [1, 5] are our baselines. For ease of exposition, we term these as RANDOM and ENGLISH-RANDOM. The former uses a concatenation of random tokens as the fingerprint key ($x_{\mathrm{fp}}$), while the latter uses a coherent English sentence sampled from Llama-3.1-8B-Instruct. For these schemes, the response ($y_{\mathrm{fp}}$) is a random unrelated token. These have been described as Random Questions and Natural Questions, resp. in prior work [1, 5].
@@ 4.1 Scalability: How many fingerprints can we add?
Scaling to a large number of fingerprints is crucial for making model sharing secure, e.g., as we show in Fig.˜6. However, existing works embed only up to 100 fingerprints [5] because ENGLISH-RANDOM fingerprint generation–English keys and random responses–suffers from significant utility drop after 256 fingerprints as seen in Fig.˜3 (left). Another baseline scheme of RANDOM–which uses a sequence of random tokens as key and response–is Scalable but not secure, because such keys can easily be detected and filtered out by model hosts at inference time. Our proposed scheme of using Perinucleus fingerprints with English keys achieves the best of both worlds – it has In-distribution keys and better Harmlessness by trading off modestly on Uniqueness (defined in Section˜3). We can hence embed 24,576 fingerprints without significant drop in model performance as seen in the plot – two orders of magnitude improvement over the existing baseline of ENGLISH-RANDOM [1, 5]. Further, as we show in App E.5, our lower Uniqueness does not lead to a high false positive rate.
Generalizability of our scheme. We demonstrate our scheme’s Scalability on various model sizes of Llama-3.1, as well as base and instruct versions of various model families [30, 29, 32, 31] (totalling to 10 models) in Fig.˜4. We find that the relative drop in performance is less than 5% across the models considered even at 8192 fingerprints. While instruct models are generally more sensitive to adding fingerprints at larger scales, Perinucleus sampling improves significantly over baselines, which can induce non-trivial performance drops at just 256 fingerprints (see Fig.˜13 in Section˜F.5), demonstrating the broader applicability of our method.
Figure 4: Performance across models. We plot the avg scores of models fingerprinted with our scheme on OpenLLM for different sized Llama 3.1 models (left) and for base (middle) and instruction-tuned (right) models from other families. We find that the relative performance is over 95% even at 8192 fingerprints across models. The x-axis is logarithmic. See Fig.˜13 for comparison to baselines.
@@ 4.2 Persistence: How many fingerprints survive SFT?
An important property of fingerprints is their ability to Persist after SFT on other data. We investigate this Persistence in Fig.˜3 (right)after 2 epochs of SFT on Alpaca [39] for a Llama-3.1-8B model.
@@
The baseline of ENGLISH-RANDOM from [1, 5] leads to fingerprints that are easily forgotten, while using RANDOM strings as keys results in higher Persistence. Since RANDOM keys are out-of-distribution from the SFT data, we posit that the changes induced by SFT do not change the model’s behavior much on RANDOM fingerprints. This leads to higher Persistence.
@@
The Perinucleus scheme also demonstrates high Persistence, retaining over 60% of fingerprints from an initial set of 8192. We hypothesize that the in-distribution nature of the responses (as compared to ENGLISH-RANDOM) leads to better Persistence. Note that Persistence decreases as more fingerprints are inserted. As the number of fingerprints increases, the average value of $p_{\theta_{\text{fp}}^{m}}(y_{\mathrm{fp}}|x_{\mathrm{fp}})$ after fingerprinting goes down (as we show in Section˜F.2), since we regularize the model to have a high utility. This means that a greater fraction of fingerprints are closer to the margin of being forgotten as we increase the number of fingerprints, and this leads to a lower Persistence. This effect is even more pronounced for schemes where $p_{\theta^{m}}(y_{\mathrm{fp}}|x_{\mathrm{fp}})$ is already low, i.e. where the response was chosen randomly (e.g. the scheme from [5]). However, the rate of this decrease is sublinear for Perinucleus fingerprints, indicating that the total number of retained fingerprints still increases as the number of fingerprints inserted is increased. We explicitly show this in Fig.˜14 in the Appendix.
As we show in our ablation study (Fig.˜12 in Appendix˜F), regularization improves the Harmlessness of all fingerprint schemes, however, better fingerprint design improves both Persistence and Harmlessness. Further, one can trade-off between the two with different regularization parameters, and we choose the operating point with the highest model utility.
How do post-training choices affect persistence? In Fig.˜5, we analyze how different post-training choices affect the persistence of Perinucleus fingerprints on a Llama-3.1-8B model on a single seed. In the plot on the left, we show the persistence after fine-tuning on a fraction of the Alpaca dataset, and find that persistence drops almost log-linearly with the number of samples. We also investigate the relationship of persistence with number of SFT epochs (Fig.˜5 middle), and find that it drops a bit before stabilizing after 2 epochs of SFT on Alpaca. Finally, we analyze the effect of the SFT dataset on persistence (Fig.˜5 right). We measure persistence after SFT on MathInstruct[40], a larger math dataset, and find that it leads to less forgetting as compared to Alpaca. We hypothesize that this happens because its prompts are farther from the fingerprints’ distribution, leading to lower interference on the model’s behavior on fingerprint keys. We also check persistence after SFT on Alpaca followed by 1 epoch of DPO [41] on Orca pairs [42], and find that this does not induce much more forgetting beyond that induced by the SFT stage, demonstrating the scheme’s robustness.
Figure 5: Effect of number of samples, epochs and dataset for fine-tuning on persistence for Llama-3.1-8B: (Left) Persistence decreases roughly log-linearly with number of SFT samples. (Middle) Persistence decreases slightly before stabilizing with increasing number of SFT epochs. (Right) Persistence is also affected by the distribution of the SFT data, with chat like data having a higher effect than Math data. Finally, additional DPO after instruction tuning does not lead to many more fingerprints being forgotten. These trends are consistent for 1024 and 4096 fingerprints.
5 Security Threats through Collusion and a Novel Defense via Scalability
Existing fingerprinting techniques of [1, 5, 15] all suffer from vulnerability against changes to how the model is used. In Appendix˜E, we address several security and robustness risks, including different sampling algorithms, merging fingerprinted and non-fingerprinted models, prompt-based attacks and false positive detection. We empirically characterize the tradeoffs involved and propose some mitigations for such risks. Scalability also provides a layer of defense against existing attacks, since it provides a higher number of fingerprints for the owner to check a suspicious model with. In this section, we introduce and focus on an under-studied threat of a collusion attack, and provide a provable defense; this exemplifies why Scalability is critical for Security.
@@ 5 Security Threats through Collusion and a Novel Defense via Scalability
One of the benefits of fingerprinting is the ability to share a model with a larger community. A natural scenario is when a model owner receives a request to share the model weights and sends a fingerprinted version of the model to a model host, who then runs some service using the model. Fingerprinting helps detect when the model is illegally copied and hosted by others without legitimate access. When another model host requests access, another copy of the model with potentially different set of fingerprints is shared, so that we can uniquely link each model with the corresponding host.
@@
Threat model. When $N$ versions of a base model are shared with $N$ model hosts, a coalition of adversarial hosts may pool their models to avoid detection. If all fingerprints are unique, i.e., no two models share any fingerprints, then such a coalition can identify and avoid answering fingerprint queries strategically. By running multiple models for each query, they can identify differences in fingerprinting because their models will respond differently. They can respond to queries using strategies to evade detection, including the following: $(i)$ Majority voting: The coalition responds with the output produced by the most models, breaking ties randomly; $(ii)$ Minority voting: The coalition responds with the output produced by the fewest models, breaking ties randomly; and $(iii)$ Non-unanimous refusal: The coalition refuses to respond to any query where there is disagreement among the models. Another flavor of collusion through model merging is studied in Appendix˜E.
@@
Novel collusion resistant fingerprinting strategy. We introduce a simple and efficient scheme to assign fingerprints and identify models (in Definition˜5.1). In Fig.˜6, we empirically demonstrate that this strategy is secure against the three standard collusion attacks and an additional Optimal attack, which we outline in the proof of Proposition˜5.3. While the Optimal strategy helps adversaries avoid detection most effectively, we can still ensure accurate detection with enough fingerprints. Together with our theoretical guarantee against all collusion attacks in Proposition˜5.3, this shows that embedding enough number of fingerprints in each model, i.e., Scalability, is critical in achieving security, i.e., identifying at least one colluding model.
@@
Note that if the coalition of attackers can respond with any other model than the fingerprinted models then it is impossible to detect the collusion. The attacker can simply choose to answer with the other model all the time, in which case the attacker is not using the fingerprinted model at all. To disallow such degenerate scenarios, we need a mild assumption in our analysis.
Assumption 5.2 (Response under unanimous output).
If all models in the coalition produce the same output, the coalition must respond accordingly.
This guarantees the detection of a single model from the coalition. In general, it is impossible to guarantee the detection of the entire coalition without stronger assumptions, because, for example, the coalition can choose to use only the responses of a single model.
@@
In the presence of collusion, the main idea is that there will be enough agreements among the coalition such that at least one of the colluding models will have a high enough score. This ensures that a large enough number of fingerprints guarantees identification.
@@
Under ˜5.2 and the fingerprinting scheme of Definition˜5.1, when there are $N$ models and a maximum coalition size of $K$, for any $\delta\in(0,1)$, there exists $p\in(0,1)$ such that
@@
fingerprints will guarantee detection of at least one model from the coalition with probability $1-\delta$.
We defer the proof to Appendix˜B. Although the bound on the number of required fingerprints scales poorly in $K$, this is unlikely to be an issue in practice because forming a coalition of size $K$ makes inference $K$ times more expensive. Thus, collusion will only be economically viable for small $K$. In contrast, the logarithmic scaling in $N$ ensures that we can support a large number of models. In Fig.˜6 on the right, we show how well our defense works quantitatively. For $N=2048$ models, under various 3-way collusion attacks, the proposed collusion resistant fingerprinting with $p=0.243$ achieves near-perfect detection rate when the number of total fingerprints $M$ is larger than $2048$. This implies that each model needs to include at least $Mp=500$ fingerprints on average to achieve security against collusion attacks. This underscores the necessity of a Scalable fingerprinting scheme.
Figure 6: Detection rate under 3-way collusion attacks.